Find 65.4 ÷ 6 Find 65.4 ÷ 6 Find 65.4 ÷ 6 Find 65.4 ÷ 6

Find 
 65.4 ÷ 6
Find 
 65.4 ÷ 6
Find 
 65.4 ÷ 6
Find 
 65.4 ÷ 6

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have,
= (654/10) ÷ 6
Then,
= (654/10) × (1/6)
= (109/10) × (1/1)
= (109/10)
On dividing a decimal by 10, the decimal point is shifted to the left by one place.
Then,
= 10.9

Thanks